OpenStudy (cherry17ann):

When 40 is added to the number of miles Karen ran last week, the result is the same as adding 10 to 4 times the number of miles she ran last week. How many miles did Karen run last week?

Directrix (directrix):

@cherry17ann Do you still want to do this problem?

OpenStudy (cherry17ann):

Sure! @directrix I got it wrong anyways

Directrix (directrix):

Okay. Let M = the number of miles Karen ran last week

Directrix (directrix):

Translate this to symbols: 40 is added to the number of miles Karen ran last week 40 + M

Directrix (directrix):

Translate this to symbols: adding 10 to 4 times the number of miles she ran last week. 10 + 4*M

Directrix (directrix):

That gives the equation: 40 + M = 10 + 4M M = ? @cherry17ann Solve for M. Post what you get, okay?

OpenStudy (cherry17ann):